Your bank will lend you $2,000 for 60 days at a cost of $35 interest. What is your effective rate of interest? Al Pay Day Lenders will lend $200 for 10 days at a cost of S 10 interest. What is the effective rate of interest?
Answer #1
Interest rate for 60 Days 1.75% =35/2000
Effective annual interest rate 110.97% =(1+1.75%)^(360/60)
Interest rate for 10 Days 0.50% =10/2000
Effective annual interest rate 119.67% =(1+0.5%)^(360/10)